OpenStudy (anonymous):

Rewrite in simplest rational exponent form √x • 4√x. Show each step of your process.

jhonyy9 (jhonyy9):

sqrtx*4sqrtx = x^(1/2) *4*x^(1/2) = 4*x^((1/2)+(1/2)) = 4*x^(2/2) = 4x

OpenStudy (anonymous):

How exactly do I read that?

OpenStudy (anonymous):

since you can multiply in any order, imagine 4 • (√x • √x) (√x • √x) = x so = 4x